OpenStudy (anonymous):

Simplify the following ratio: 2.8m : 70cm

OpenStudy (anonymous):

???

OpenStudy (anonymous):

4:1

OpenStudy (anonymous):

2.8m:0.7m =2.8/0.7 =4/1 =4:1

OpenStudy (anonymous):

\[\frac{280cm}{70cm}=\frac{4}{1}\] Multiply m by 100 to convert to cm

OpenStudy (anonymous):

Thx :D